4. A combustion reaction wherein 15.0 g of propane (C3H8) is burned with 75.0 g of oxygen (O2).


a. Determine the limiting reactant.


b. What is the maximum amount of CO2 (g) can you produce?.



1
Expert's answer
2022-01-17T13:50:04-0500

Solution:

The molar mass of C3H8 is 44.1 g/mol

The molar mass of O2 is 32.0 g/mol


Calculate the number of moles of each reactant:

(15.0 g C3H8) × (1 mol C3H8 / 44.1 g C3H8) = 0.34 mol C3H8

(75.0 g O2) × (1 mol O2 / 32.0 g O2) = 2.34 mol O2


Balanced chemical equation:

C3H8(g) + 5O2(g) ⟶ 3CO2(g) + 4H2O(g)

According to stoichiometry:

1 mol of C3H8 reacts with 5 mol of O2

Thus, 0.34 mol of C3H8 reacts with:

(0.34 mol C3H8) × (5 mol O2 / 1 mol C3H8) = 1.7 mol O2

However, initially there is 2.34 mol of O2 (according to the task).

Thus, C3H8 acts as limiting reactant and O2 is excess reactant.


According to stoichiometry:

1 mol of C3H8 produces 3 mol of CO2

Thus, 0.34 mol of C3H8 produces:

(0.34 mol C3H8) × (3 mol CO2 / 1 mol C3H8) = 1.02 mol CO2


The molar mass of CO2 is 44.01 g/mol

Hence,

(1.02 mol CO2) × (44.01 g CO2 / 1 mol CO2) = 44.89 g CO2


Answer:

a) The limiting reactant is propane (C3H8)

b) The maximum amount of CO2 is 1.02 mol (or 44.89 g)